Please help me- What is the slope of the line y=2/3x+2
Answer by stanbon(75887)

What is the slope of the line y=2/3x+2
-----
slope is the change in y when x changes 1.
----
Let x = 0, then y = (2/3)0+2 = 2
---
Let x = 1, then y = (2/3)1 + 2 = 2+(2/3)
--------
Notice that y has increased by 2/3 when x increased by 1.
---
So the slope is 2/3
========================
Cheers,
Stan H.